# RAGONITE: Iterative Retrieval on Induced Databases and Verbalized RDF for Conversational QA over KGs with RAG

### v2

@@ 3 Demonstration walkthrough
Walkthrough. We go through only the main workflow here: everything else can be explored by a user during an interactive demo session. Offline, the backend KG is preprocessed to derive the equivalent DB and the text corpus. At runtime, in Step 1, users enter their conversational question (potentially intent-implicit) in the input box provided. Suppose we are at the second conversation turn now (Sounds like these are sleek sedans. Is an average X1 taller than a Coupe?). In Step 2a, this question is passed on to the intent-explicit SQL-formulating LLM call (Sec. 2.1). In this case, the LM decides that results of one iteration (that only retrieved heights of BMW X1 variants) was not enough to satisfy the intent in this complex question (Sec. 2.3), and so it formulates a second query (this time fetching the heights of Coupe models). At this moment, the iterator stops SQL querying and invokes the text retriever (Step 2b, Sec. 2.2) via an intent-explicit NL question (BMW X1 average height compared to BMW Coupe Models). Note that Steps 2a and 2b can happen in an arbitrary order. A sample verbalization of a Turtle capsule can be seen in the screenshot below the corresponding blue box (for the entity “bmw x1 xdrive 25e sport”, we lowercase all text for simplicity), along with the retrieval score assigned by the vector search model (0.01700170.0170.017). The iterator decides that contents of one round of text retrieval has enough information pertinent to the question, and by now both tools have been explored, so it hands over these results to the answer generator (Sec. 2.4) This LLM agent crafts the final response shown to the user (Step 3). Through explicit prompting, we make the LLM inject source citations in its answer for scrutability. In this case, sources [1-2] refer to SQL results and [3-7] to verbalizations (totaling $10$ sources, where $10$ is configurable). Prompts for all LLM calls can be examined by an user via the answer derivation panel.
Backend. The core of RAGonite’s backend handles query formulation, SQL execution, text retrieval, iteration/tool selection, and answer generation. The backend also has a stateful layer that stores conversations in a SQLite DB and provides a REST API to the frontend with FastAPI. Core dependencies include the vector database (ChromaDB), a template engine used for prompts (Jinja), and LLM libraries (OpenAI and Ollama). ChromaDB is used as our vector DB for storing RDF verbalizations. We also use ChromaDB’s in-built vector search. For all of our experiments, we ran RAGONITE on a GPU server (4x48GB NVIDIA Ada 6000 RTX, 512 GB RAM, 64 virtual cores) which is shared with other researchers. A single GPU was dedicated to running Llama-3.3-70B. If the LLM used for text inference is hosted in the cloud, for example, by using an OpenAI model, the remaining resource requirements are minimal and the RAGonite instance can also run on a standard laptop.
@@ 4 Evaluation
Results. The two-pronged iterative approach did the best ($28/30$ questions correct). SQL-only got $18$ (failure to handle abstract intents) and verbalization-only got $24$ cases right (failure to handle complex math). A baseline where GPT-4o generated SPARQL queries, only got $4$ answers correct. A particular vulnerability of SQL and SPARQL was incorrect entity linking for ad hoc mentions (like “gran coupe sport 220 i m” in questions, while the exact entity label is bmw-220i-m-sport-gran-coupé), and this is where verbalizations are particularly helpful. This also proved that LLMs are still much weaker at SPARQL generation than SQL (presumably due to lower training data), especially when it comes to complex or abstract intents in ConvQA.
@@ 5 Conclusions and future work
We showcase RAGONITE, a transparent RAG pipeline that adopts a novel alternative approach to ConvQA over KGs by iterative retrieval and merging results of structured and unstructured querying. While Text2SPARQL has been investigated in the past, we adopt a novel alternative of exploiting Text2SQL, where LLMs do much better, over a DB that is programmatically induced from the KG. By converting a KG to its verbalized form that is indexed as a vector DB, we show that the information in knowledge graphs can be harnessed for more abstract intents as well. RAGONITE contains an agentic workflow: key future work would enhance this further by incorporating reflection mechanisms that enable each module to critique and improve its output, without compromising efficiency. Another direction would be to fine-tune specific pipeline components for a RAG setting with synthetic data.

### v3

@@ 3 Demonstration walkthrough
Walkthrough. We go through only the main workflow here: everything else can be explored by a user during an interactive demo session. Offline, the backend KG is preprocessed to derive the equivalent DB and the text corpus. At runtime, in Step 1, users enter their conversational question (potentially intent-implicit) in the input box provided. Suppose we are at the second conversation turn now (Sounds like these are sleek sedans. Is an average X1 taller than a Coupe?). In Step 2a, this question is passed on to the intent-explicit SQL-formulating LLM call (Sec. 2.1). In this case, the LLM decides that results of one iteration (that only retrieved heights of BMW X1 variants) were not enough to satisfy the intent in this complex question (Sec. 2.3), and so it formulates a second query (this time fetching the heights of Coupe models). At this moment, the iterator stops SQL querying and invokes the text retriever (Step 2b, Sec. 2.2) via an intent-explicit NL question (BMW X1 average height compared to BMW Coupe Models). Note that Steps 2a and 2b can happen in an arbitrary order. A sample verbalization of a Turtle capsule can be seen in the screenshot below the corresponding blue box (for the entity bmw-x1-xdrive-25e-sport, we lowercase all text for simplicity), along with the retrieval score assigned by the vector search model (0.01700170.0170.017). The iterator decides that contents of one round of text retrieval has enough information pertinent to the question, and by now both tools have been explored, so it hands over these results to the answer generator (Sec. 2.4) This LLM agent crafts the final response shown to the user (Step 3). Through explicit prompting, we make the LLM inject source citations in its answer for scrutability. In this case, sources [1-2] refer to SQL results and [3-7] to verbalizations (understandable by labels above respective results). Prompts for all LLM calls can be examined by an user via the answer derivation panel.
Backend. The core of RAGonite’s backend handles query formulation, SQL execution, text retrieval, iteration/tool selection, and answer generation. The backend also has a stateful layer that stores conversations in a SQLite DB and provides a REST API to the frontend with FastAPI. Core dependencies include the vector database (ChromaDB), a template engine used for prompts (Jinja), and LLM libraries (OpenAI and Ollama). ChromaDB is used as our vector DB for storing RDF verbalizations. We also use ChromaDB’s in-built vector search. For all of our experiments, we ran RAGONITE on a shared GPU server (4x48GB NVIDIA Ada 6000 RTX, 512 GB RAM, 64 virtual cores). A single GPU was dedicated to running Llama-3.3-70B. If the LLM used for text inference is hosted in the cloud, for example, by using an OpenAI model, the remaining resource requirements are minimal and the RAGonite instance can also run on a standard laptop.
@@ 4 Evaluation
Results. The two-pronged iterative approach did the best ($28/30$ questions correct). SQL-only got $18$ (failure to handle abstract intents) and verbalization-only got $24$ cases right (failure to handle complex math). A baseline where GPT-4o generated SPARQL queries, got only $4$ answers correct. A particular vulnerability of SQL and SPARQL was incorrect entity linking for ad hoc mentions (like “gran coupe sport 220 i m” in questions, while the exact entity label is bmw-220i-m-sport-gran-coupé), and this is where verbalizations are particularly helpful. This also proved that LLMs are still much weaker at SPARQL generation than SQL (presumably due to lower training data), especially when it comes to complex or abstract intents in conversational question answering.
@@ 5 Conclusions and future work
We showcase RAGONITE, a transparent RAG pipeline that adopts a novel alternative approach to ConvQA over KGs by iterative retrieval and merging results of structured and unstructured querying. While Text2SPARQL has been investigated in the past, we adopt a novel alternative of exploiting Text2SQL, where LLMs do much better, over a DB that is programmatically induced from the KG. By converting a KG to its verbalized form that is indexed as a vector DB, we show that the information in knowledge graphs can be useful for more abstract intents as well. RAGONITE contains an agentic workflow: key future work would enhance this further by incorporating reflection mechanisms that enable each module to critique and improve its output, without compromising efficiency. Another direction would be to fine-tune specific pipeline components for a RAG setting with synthetic data.